# TRG-Net: An Interpretable and Controllable Rain Generator

### v1

@@ 4 Transformable Rain Generator
where $\mathcal{X}$, $\mathcal{R}$, $\mathcal{B}\in\mathbb{R}^{H\times W\times 3}$ are RGB images, denoting the rainy image, the rain layer image, and the clean background image, respectively. $\tilde{\mathcal{C}}\in\mathbb{R}^{p\times p\times 3\times K}$ is a rain kernel tensor, representing $K$ rain kernels, ~$\tilde{\mathcal{M}}\in\mathbb{R}^{H\times W\times K}$ denotes $K$ rain maps, $H$ and $W$ represent the height and width of the images, respectively, and $p$ denotes the spatial size of the rain kernel. Although $\tilde{\mathcal{C}}$ and $\tilde{\mathcal{M}}$ can be learned or generated by deep learning based manners, they are all fixed after the model is trained, which neglects essential dynamic rain factors like shape, orientation, length and so on. Therefore, the model (9) is actually only suitable in deraining tasks, but not suitable for generating rainy images with diverse rain factors. Besides, this model simply adds the rain layer to the background image, which is also just a rough approximation to the real complex situations with fog-like rains.
@@
4.1 Proposed Backbone Rain Model
@@
$\begin{split}&\mathcal{X}(\theta,s_{l},s_{w},\bm{\alpha},\tau)\\ &~{}~{}~{}~{}~{}~{}=\mbox{MerNet}\left(\mathcal{R}(\theta,s_{l},s_{w},\bm{\alpha},\tau),\mathcal{B}\right)\\ &~{}~{}~{}~{}~{}~{}=\mbox{MerNet}\left(\mathcal{C}(\theta,s_{l},s_{w},\bm{\alpha})\otimes\mathcal{M}(\tau),\mathcal{B}\right),\end{split}$ (10)
where $\theta$, $s_{l}$, $s_{w}$, $\bm{\alpha}$ and $\tau$ represent the orientation, length, width, shape, and sparsity parameters, respectively. $\mathcal{X}(\theta,s_{l},s_{w},\bm{\alpha},\tau)$, $\mathcal{R}(\theta,s_{l},s_{w},\bm{\alpha},\tau)$ and $\mathcal{B}$ denote the parameterized rainy image, the parameterized rain layer and the input clean background image, respectively. $\mathcal{C}(\theta,s_{l},s_{w},\bm{\alpha})\in\mathbb{R}^{p\times p\times 3\times K}$ and $\mathcal{M}(\tau)\in\mathbb{R}^{H\times W\times K}$ are the rain kernel and the rain map, corresponding to $\tilde{\mathcal{C}}$ and $\tilde{\mathcal{M}}$ in Eq. (9), respectively. $\mbox{MerNet}(\cdot)$ is a deep network for merging the rain layer and background image, which simulates the interaction between rain and background image, such as the rain streak photometry and fog-like rain.
Figure 4: The proposed transformable rainy image generation network (TRG-Net), which takes the proposed rain model (10) as the backbone. Similar to the conventional artificial rainy image synthesis framework as shown in Fig. 1, it also consists of three parts: (a) the rain kernel model (Sec. 4.2.1), (b) the rain map model (Sec. 4.2.2) and (c) the merging model (Sec. 4.2.3).
@@
4.2 Transformable Rainy Image Generator
@@
4.2.1 Rain Kernel Generation
@@
where $\mathcal{D}(\theta,s_{l},s_{w})\in\mathbb{R}^{p\times p\times 3\times M}$ represents the rain kernel dictionary with $M$ elements. $\bm{\alpha}\in\mathbb{R}^{M\times K}$ denotes the $M$-dimensional coefficient vectors for the $K$ rain kernels, $\odot$ is the tensor multiplication along the $4^{th}$ dimension [^4] For $\mathcal{D}\in\mathbb{R}^{p\times p\times 3\times M}$ and $\bm{\alpha}\in\mathbb{R}^{M\times K}$, $(\mathcal{D}\odot\bm{\alpha})\in\mathbb{R}^{p\times p\times 3\times K}$ and $(\mathcal{D}\odot\bm{\alpha})_{\{:,:,:,k\}}=\sum_{m=1}^{M}\mathcal{D}_{\{:,:,:,m\}}\bm{\alpha}_{mk}$.. Besides, $\mathcal{D}(\theta,s_{l},s_{w})$ is parameterized with the manner in Eq. (5). By substituting Eqs. (5) and (8), we can obtain that for any $i,j=1,2,\cdots,p$, $c=1,2,3$ and $k=1,2,\cdots,K$,
$\begin{split}\!\left[\mathcal{C}(\theta,\!s_{l},\!s_{w},\!\bm{\alpha})\right]_{ijck}\!\!\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\ =\!\sum_{mn}\!\bm{\alpha}_{mk}\!\cdot\!w_{mnc}^{d}\!\cdot\!\varphi_{n}\!\left(T_{\{\theta,s_{l},s_{w}\}}\!\cdot\!x_{ij}\right),\end{split}$ (12)
@@
Rain kernel generator. It is easy to find that the calculation about $\theta,s_{l},s_{w}$ and $\bm{\alpha}$ only involves simple operations, all of which can be easily performed using popular deep learning tools, such as Pytorch, Tensorflow, and general gradient backpropagation algorithms. Therefore, we can adopt lightweight fully connection networks (FCNs) for learning the distribution of these parameters, and train these FCNs with other parts of the proposed rainy image generator in an end-to-end way. Specifically, as shown in Fig. 4 (a), we exploit the following four FCNs for rain factors generation [^5] There is a reshape in output of the FCN for $\bm{\alpha}$.:
@@
where all FCNs are input with Gaussian noise, and $\bm{w}^{\theta}$, $\bm{w}^{l}$, $\bm{w}^{w}$ and $\bm{w}^{\alpha}$ denote the to-be-learned parameters of the FCNs. Combining Eqs. (12) and (13), the entire rain kennel generator is
$\begin{split}&\mathcal{C}_{ijck}=\sum_{mn}\!w_{mnc}^{d}\cdot\left[\mbox{FCN}^{\alpha}_{\bm{w}^{\alpha}}(e_{\alpha})\right]_{mk}\cdot\\ &\varphi_{n}\!\left(\!T_{\left\{\mbox{\scriptsize{FCN}}^{\theta}_{\bm{w}^{\theta}}(e_{\theta}),\mbox{\scriptsize{FCN}}^{l}_{\bm{w}^{l}}(e_{l}),\mbox{\scriptsize{FCN}}^{w}_{\bm{w}^{w}}(e_{w})\right\}}\!x_{ij}\!\right),\!\end{split}$ (14)
@@ 4.2.2 Rain map Generation
Rain map generator. Similar Eq. (13) for learning other rain factors, we adopt a lightweight FCN to learn the distribution of sparsity factor $\tau$:
@@
$\begin{split}\mathcal{M}=\mbox{ReLU}\Big{(}&\mbox{rotResNet}_{\bm{w}^{r}}\!\left(\!\mathcal{Z},\mbox{FCN}^{\theta}_{\bm{w}^{\theta}}(e_{\theta})\!\right)\Big{.}\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\Big{.}-\mbox{FCN}^{\tau}_{\bm{w}^{\tau}}(e_{\tau})\Big{)},\end{split}$ (19)
@@
where $\bm{w}^{m}=\left\{\bm{w}^{r},\bm{w}^{\tau},\bm{w}^{\theta}\right\}$ represents all the to-be-learned parameters.
4.2.3 Merging Rain Layer and Background
@@
$\begin{split}&\mathcal{X}(\theta,s_{l},s_{w},\bm{\alpha},\tau)=\\ &~{}~{}\mbox{rotResNet}_{\bm{w}^{o}}\left(\mbox{cat}(\mathcal{R}(\theta,s_{l},s_{w},\bm{\alpha},\tau),\mathcal{B}),\theta\right),\!\\ \end{split}$ (21)
@@
where $\bm{w}^{x}=\{\bm{w}^{o},\bm{w}^{\theta}\}$ denotes the to-be-learnt parameters for merging network.
@@
$\begin{split}\mathcal{X}&=\mbox{TRG-Net}_{\bm{w}}\left(\mathcal{B},\mathcal{Z},\bm{e}\right)\\ &\triangleq\mbox{MerNet}_{\bm{w}^{x}}\left(\mbox{KerNet}_{\bm{w}^{c}}(e_{\theta},e_{l},e_{w},e_{\alpha})\otimes\right.\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\mbox{MapNet}_{\bm{w}^{m}}(\mathcal{Z},e_{\tau},e_{\theta}),\mathcal{B},e_{\theta}\right).\end{split}$ (24)
@@
Remark. It should be noted that the proposed framework not only possesses intrinsic interpretability, where the distribution of rain factors can be purely learned from data using GAN-based manners, but also has essential controllability. Specifically, after achieving $\bm{w}$ through learning from data, we can obtain all parameters. Then either manually-set rain factors or FCN-generated rain factors can be utilized as the inputs to model (21). When using manually-set rain factors, the generated rainy images are essentially controlled. These merits are finely validated in Fig. 7 below.
4.3 Loss Function
@@
Fortunately, in the proposed method, the orientation degree $\theta$ of the generated rain layer is able to be readily extracted by our model without extra effort. Moreover, the proposed filter parameterization and transformable convolution techniques allow us to design a rotated differential filter. By adopting the transformable kernel (5) and the inverse rotation matrix (7), the proposed rotated differential filter can be naturally defined. For $\forall i,j=1,2,\cdots,p$, we have
$\begin{split}[D(\theta)]_{ij}=\sum_{n=1}^{N}w_{n}\varphi_{n}(T_{\{\theta\}}\cdot x_{ij}),\end{split}$ (27)
Then, we can estimate $w_{n}$ by solving $D(0)=\hat{D}$, where $\hat{D}$ is the vertical difference filter. That is, we need to solve the following linear equation:
$\sum_{n=1}^{N}w_{n}\varphi_{n}(x_{ij})=\hat{D}_{ij},\forall i,j=1,2,\cdots,p,$ (28)
Actually, the solution to this equation is quite simple. Since FSE-FP used in our model is equivalent to the discrete inverse Fourier transform when $\theta=0$, $w_{n}$ can be solved by adopting discrete Fourier transform on $\hat{D}_{ij}$, i.e., $w_{n}=\nicefrac{{1}}{{p^{2}}}\sum_{st}\hat{D}_{st}\varphi_{n}(x_{st})$. Then $\forall i,j=1,2,\cdots,p$, we have
@@
Then the proposed rotatable total variation (rotTV) regularizer can be explicitly calculated by
@@
4.4 Implementation Details
@@
Our experiments empirically show that a single convolution of the rain kernel and the rain map is usually not sufficient to generate the diverse, complex and long rain streaks. Therefore, we actually generate three rain kernels and convolute them with the rain maps to increase the receptive field of rain kernel.
Besides, it is easy to find that the rain kernels contain a clear physical meaning, which is quite different from other convolution kernels in deep networks. Therefore, we specifically design an initialization manner for them, which achieves a streak-like initialization. Please refer to the appendix for more details for the rain kernels initialization and generation.
@@ 5 Experimental Results
5.1 Rain Generator Verification
The proposed rain generator learns the rain factors by FCNs from data samples. This allows us to extract their distributions purely from the rain data set. We thus verify this capability on Rain100L [27], a synthetic rain dataset containing 200 pairs of rainy and rainless images for training.
Experiment settings. To train TRG-Net by an adversarial strategy, following [22] and [35], we adopt a patch-based discriminator [17]. Following [18], the initial learning rates for the generator and the discriminator are $1\times 10^{-4}$ and $4\times 10^{-4}$, respectively. As suggested in [47], we update the discriminator five times for each generator updating. We utilize Adam algorithm [48] to optimize the rain generation network. We train the proposed generator for 200 epochs with rotatable TV regularizer, and the network is trained for $3000$ iterations in each epoch. We randomly crop a $256\times 256$ patch as input in each iteration.
The learned distributions [^6] The distribution is approximated by the histogram obtained by sampling 10000 times. of rain factors in Rain100L are shown in Fig. 7, which can be easily used to further generate different rain types by adjusting rain factors in a controllable manner. As shown in Fig. 7 (a), the learned orientation degree $\theta$ distribution is close to a Gaussian distribution, with values mostly between $-40$ and $40$, which complies with the intuitive orientation distribution in Rain100L [5]. By model (21), we can manually input the orientation degree into the trained generator to control the orientation of the generated rain steaks. Fig. 7 (a) shows the rain generated with input $\theta$ of $0^{\circ}$, $30^{\circ}$, $60^{\circ}$, $-30^{\circ}$, respectively. It should be noted that this manner offers the capability to generate OOD rains, e.g., the $60^{\circ}$ rain in Fig. 7 (a), which will be beneficial for deraining on OOD rain data sets (see OOD rain augmentation in Sec. 5.3.3). Yet this is difficult to be achieved by traditional DL-based generators, since black-box model can hardly generalize samples beyond the distribution of training data. Similarly, in Fig. 7 (b), (c) and (d), we can observe that the proposed method can also learn the distributions of $s_{l}$, $s_{w}$ and $\tau$, respectively. We can control rain from thin to thick and from short to long by scaling the rain kernels according to different $s_{w}$ and $s_{l}$ values, respectively, and control the sparsity of rain by adjusting the input of $\tau$.

### v2

@@ IV Transformable Rain Generator
where $\mathcal{X}$, $\mathcal{R}$, $\mathcal{B}\in\mathbb{R}^{H\times W\times 3}$ are RGB images, denoting the rainy image, the rain layer image, and the clean background image, respectively. $\tilde{\mathcal{C}}\in\mathbb{R}^{p\times p\times 3\times K}$ is a rain kernel tensor, representing $K$ rain kernels, ~$\tilde{\mathcal{M}}\in\mathbb{R}^{H\times W\times K}$ denotes $K$ rain maps, $H$ and $W$ represent the height and width of the images, respectively, and $p$ denotes the spatial size of the rain kernel. Although $\tilde{\mathcal{C}}$ and $\tilde{\mathcal{M}}$ can be learned or generated by deep learning based manners, they are all fixed after the model is trained, which neglects essential dynamic rain factors like shape, orientation, length and so on. Therefore, the model (9) is actually only suitable in deraining tasks, but not suitable for generating rainy images with diverse rain factors. Besides, this model simply adds the rain layer to the background image, which is also just a rough approximation to the real more complex situations.
@@
IV-A Proposed Backbone Rain Model
@@
$\begin{split}\!\mathcal{X}\!(\theta,s_{l},s_{w},\bm{\alpha},\tau)&=\!\mbox{MerNet}\left(\mathcal{R}(\theta,s_{l},s_{w},\bm{\alpha},\tau),\mathcal{B}\right)\\ &=\!\mbox{MerNet}\!\left(\!\mathcal{C}(\theta,s_{l},s_{w},\bm{\alpha})\!\otimes\!\mathcal{M}(\tau),\mathcal{B}\right),\end{split}$ (10)
where $\theta$, $s_{l}$, $s_{w}$, $\bm{\alpha}$ and $\tau$ represent the orientation, length, width, shape, and sparsity parameters, respectively. $\mathcal{X}(\theta,s_{l},s_{w},\bm{\alpha},\tau)$, $\mathcal{R}(\theta,s_{l},s_{w},\bm{\alpha},\tau)$ and $\mathcal{B}$ denote the parameterized rainy image, the parameterized rain layer and the input clean background image, respectively. $\mathcal{C}(\theta,s_{l},s_{w},\bm{\alpha})\in\mathbb{R}^{p\times p\times 3\times K}$ and $\mathcal{M}(\tau)\in\mathbb{R}^{H\times W\times K}$ are the rain kernel and the rain map, corresponding to $\tilde{\mathcal{C}}$ and $\tilde{\mathcal{M}}$ in Eq. (9), respectively. $\mbox{MerNet}(\cdot)$ is a merging network for merging the rain layer and background image, which simulates the interaction between rain and background image.
Figure 4: The proposed transformable rainy image generation network (TRG-Net), which takes the proposed rain model (10) as the backbone. Similar to the conventional artificial rainy image synthesis framework as shown in Fig. 1, it also consists of three parts: (a) the rain kernel model (Sec. IV-B1), (b) the rain map model (Sec. IV-B2) and (c) the merging model (Sec. IV-B3).
@@
IV-B Transformable Rainy Image Generator
@@
IV-B1 Rain Kernel Generation
@@
where $\mathcal{D}(\theta,s_{l},s_{w})\in\mathbb{R}^{p\times p\times 3\times M}$ represents the rain kernel dictionary with $M$ elements. $\bm{\alpha}\in\mathbb{R}^{M\times K}$ denotes the $M$-dimensional coefficient vectors for the $K$ rain kernels, $\odot$ is the tensor multiplication along the $4^{th}$ dimension [^4] For $\mathcal{D}\in\mathbb{R}^{p\times p\times 3\times M}$ and $\bm{\alpha}\in\mathbb{R}^{M\times K}$, $(\mathcal{D}\odot\bm{\alpha})\in\mathbb{R}^{p\times p\times 3\times K}$ and $(\mathcal{D}\odot\bm{\alpha})_{\{:,:,:,k\}}=\sum_{m=1}^{M}\mathcal{D}_{\{:,:,:,m\}}\bm{\alpha}_{mk}$.. Besides, $\mathcal{D}(\theta,s_{l},s_{w})$ is parameterized in the manner of Eq. (5). By substituting Eqs. (5) and (8), we can obtain that for any $i,j=1,2,\cdots,p$, $c=1,2,3$ and $k=1,2,\cdots,K$,
$\begin{split}\!\left[\mathcal{C}(\theta,\!s_{l},\!s_{w},\!\bm{\alpha})\right]_{ijck}\!=\!\sum_{mn}\!\bm{\alpha}_{mk}\!\cdot\!w_{mnc}^{d}\!\cdot\!\varphi_{n}\!\left(\!T_{\{\theta,s_{l},s_{w}\}}\!\cdot\!x_{ij}\!\right)\!,\end{split}$ (12)
@@
Rain kernel generator. It is easy to find that the calculation about $\theta,s_{l},s_{w}$ and $\bm{\alpha}$ only involves simple operations, all of which can be easily performed using popular deep learning tools, such as Pytorch, Tensorflow, and general gradient backpropagation algorithms, allowing us to learn their distribution automatically from data rather than manually set. Therefore, we can adopt lightweight fully connection networks (FCNs) for learning the distribution of these parameters, and train these FCNs with other parts of the proposed rainy image generator in an end-to-end way. Specifically, as shown in Fig. 4 (a), we exploit the following four FCNs for rain factors generation [^5] There is a reshape in output of the FCN for $\bm{\alpha}$.:
@@
where all FCNs are input with Gaussian noise, and $\bm{w}^{\theta}$, $\bm{w}^{l}$, $\bm{w}^{w}$ and $\bm{w}^{\alpha}$ denote the to-be-learned parameters of the FCNs. Combining Eqs. (12) and (13), we can obtain the entire rain kennel generator,
$\begin{split}\mathcal{C}_{ijck}=&\sum_{mn}w_{mnc}^{d}\cdot\left[\mbox{FCN}^{\alpha}_{\bm{w}^{\alpha}}(e_{\alpha})\right]_{mk}\cdot\\ &~{}~{}\varphi_{n}\!\left(\!T_{\left\{\mbox{\scriptsize{FCN}}^{\theta}_{\bm{w}^{\theta}}(e_{\theta}),\mbox{\scriptsize{FCN}}^{l}_{\bm{w}^{l}}(e_{l}),\mbox{\scriptsize{FCN}}^{w}_{\bm{w}^{w}}(e_{w})\right\}}\!\cdot\!x_{ij}\!\right),\!\end{split}$ (14)
@@ IV-B2 Rain map Generation
Rain map generator. Similar to Eq. (13) for learning other rain factors, we adopt a lightweight FCN to learn the distribution of sparsity factor $\tau$:
@@
$\begin{split}\!\mathcal{M}\!=\!\mbox{ReLU}\!\Big{(}\!\mbox{rotResNet}_{\bm{w}^{r}}\!\left(\!\mathcal{Z},\!\mbox{FCN}^{\theta}_{\bm{w}^{\theta}}\!(\!e_{\theta}\!)\!\right)\!-\!\mbox{FCN}^{\tau}_{\bm{w}^{\tau}}\!(\!e_{\tau}\!)\!\Big{)}\!,\end{split}$ (19)
@@
where $\bm{w}^{m}=\left\{\bm{w}^{r},\bm{w}^{\tau},\bm{w}^{\theta}\right\}$ represents all the to-be-learned parameters. $\bm{w}^{\theta}$ and $e_{\theta}$ are shared with Eq. (15)
IV-B3 Merging Rain Layer and Background
@@
$\begin{split}\!\mathcal{X}\!(\!\theta,\!s_{l},\!s_{w},\!\bm{\alpha},\!\tau\!)\!=\!\mbox{rotResNet}_{\bm{w}^{o}}\!\left(\!\mbox{cat}(\!\mathcal{R}(\theta,\!s_{l},\!s_{w},\!\bm{\alpha},\!\tau\!),\!\mathcal{B}),\!\theta\!\right)\!,\!\\ \end{split}$ (21)
@@
where $\bm{w}^{x}=\{\bm{w}^{o},\bm{w}^{\theta}\}$ denotes the to-be-learnt parameters.
@@
$\begin{split}\mathcal{X}&=\mbox{TRG-Net}_{\bm{w}}\left(\mathcal{B},\mathcal{Z},\bm{e}\right)\\ &\triangleq\mbox{MerNet}_{\bm{w}^{x}}\left(\mbox{KerNet}_{\bm{w}^{c}}(e_{\theta},e_{l},e_{w},e_{\alpha})\otimes\right.\\ &~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\mbox{MapNet}_{\bm{w}^{m}}(\mathcal{Z},e_{\tau},e_{\theta}),\mathcal{B},e_{\theta}\right).\end{split}$ (24)
@@
Remark. It should be noted that the proposed framework is able to extract the distribution of rain factors purely from rain data without the need for rain factor labels using GAN-based manners due to these rain factors being intrinsically embedded. Besides, the proposed model possesses essential controllability. Specifically, after achieving $\bm{w}$ through learning from data, we can obtain all parameters. Then either manually-set rain factors or FCN-generated rain factors can be utilized as the inputs to model (21). When using manually-set rain factors, the generated rainy images are essentially controlled. These merits are finely validated in Fig. 7 below.
IV-C Loss Function
@@
Fortunately, in the proposed method, the orientation degree $\theta$ of the generated rain layer is able to be readily extracted by our model without extra effort. By adopting the transformable kernel (5) and the inverse rotation matrix (7), the proposed rotated differential filter can be defined [^6] The detail of derivation can be found in supplementary material.. For $\forall i,j=1,2,\cdots,p$, we have
@@
where $\hat{D}$ is the vertical difference filter. $\theta$ denotes the rain orientation. Then the proposed rotatable total variation (rotTV) regularizer can be explicitly calculated by
@@
IV-D Implementation Details
To generate the diverse, complex and long rain streaks in real world, we actually generate three rain kernels and convolute them with the rain maps to increase the receptive field of rain kernel.
Besides, it is easy to find that the rain kernels contain a clear physical meaning, which is quite different from other convolution kernels in deep networks. Therefore, we specifically design an initialization manner for them, achieving a streak-like initialization. Please refer to the supplemental material for more details for the rain kernels initialization and generation.
@@ V Experimental Results
V-A Rain Generator Verification
The proposed rain generator learns the rain factors by FCNs from data samples, which allows us to extract their distributions purely from the rain data set. We thus verify this capability on Rain100L [45], a synthetic rain dataset containing 200 pairs of rainy and rainless images for training.
Experiment settings. To train TRG-Net by an adversarial strategy, following [23] and [33], we adopt a patch-based discriminator [17]. Following [18], the initial learning rates for the generator and the discriminator are $1\times 10^{-4}$ and $4\times 10^{-4}$, respectively. As suggested in [46], we update the discriminator five times for each generator updating. We utilize Adam algorithm to optimize the rain generation network. We train the proposed generator for 200 epochs with rotatable TV regularizer, and the network is trained for $3000$ iterations in each epoch. We randomly crop a $256\times 256$ patch as input in each iteration.